# eKalibr-Stereo: Continuous-Time Spatiotemporal Calibration for Event-Based Stereo Visual Systems

### v1

@@ IV-D Evaluation of Computation Consumption
Configuration OS Name Ubuntu 20.04.6 LTS 64-Bit
@@
Grid Tracking Optimization Total
@@ V Conclusion
In this article, we present the proposed continuous-time-based spatiotemporal calibrator for event-based stereo visual systems, named eKalibe-Stereo, which is event-only and can accurately estimate both extrinsic and temporal parameters of the sensor suite. To improve the continuity of gird tracking, building upon eKalibr, an additional efficient procedure is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked complete and incomplete grid patterns, a two-step initialization is first performed to recover the initial guesses of all parameters in the estimator, followed by a continuous-time batch optimization to refine all parameters to the optimal states. Extensive real-world experiments were conducted to evaluate the performance of the eKalibr-Stereo regarding grid tracking and spatiotemporal calibration. The results indicate that eKalibr-Stereo significantly improves the event-based grid tracking rate and could achieve spatiotemporal calibration accuracy comparable to frame-based stereo visual calibrators.
Acknowledgments
Shuolong Chen would like to express his gratitude to the School of Geodesy and Geomatics, Wuhan University, and the GREAT team for providing experimental equipment.

### v2

@@ IV-D Evaluation of Computation Consumption
Config. OS Name Ubuntu 20.04.6 LTS 64-Bit
@@
Grid Tracking (Cmp. + Incmp.) Optimization Total
@@ V Conclusion
In this article, we present the proposed continuous-time-based spatiotemporal calibrator for event-based stereo visual systems, named eKalibr-Stereo, which is event-only and can accurately estimate both extrinsic and temporal parameters of the sensor suite. To improve the continuity of grid tracking, building upon eKalibr, an additional efficient procedure is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked complete and incomplete grid patterns, a two-step initialization is first performed to recover the initial guesses of all parameters in the estimator, followed by a continuous-time batch optimization to refine all parameters to the optimal states. Extensive real-world experiments were conducted to evaluate the performance of the eKalibr-Stereo regarding grid tracking and spatiotemporal calibration. The results indicate that eKalibr-Stereo significantly improves the event-based grid tracking rate and could achieve spatiotemporal calibration accuracy comparable to frame-based stereo visual calibrators.